A metro train moves 140km/hr east for an hour and then moves 120km/hr west for an hour. describe how you will find the average velocity. What is the displacement in this problem and what is the average velocity in this problem?
Answer #1

If we assume that the unit vector along the east direction is 'i' , then

displacement of the train in its 2 hours journey, d = (140X1)i - (120X1)i =20 i km

average velocity, v = d / t [ where, t = 2 hours] = (20/2) i =10 i km/h

i.e average velocity = 10 km/h in the east direction (Ans)